OpenStudy (sydthekid913):

If given f(x) = –4x – 2 and g(x) = 2x, what is (f · g)(x)?

OpenStudy (solomonzelman):

multiply what f(x) is equal to and what f(g) is equal to together to get (f · g)(x)

OpenStudy (sydthekid913):

I got \[\left( f*g \right)\left( x \right)=-8x ^{2}-4x\]

OpenStudy (sydthekid913):

is this right?

OpenStudy (sidsiddhartha):

(f · g)(x)=f(g(x))

OpenStudy (sydthekid913):

maybe? is it?

OpenStudy (sydthekid913):

that is not one of my choices but -8x^2 is